OpenStudy (anonymous):

The perimeter of a rectangular lot is 56 meters. The length exceeds the width by 12 meters. Find the length of the lot.

OpenStudy (anonymous):

14 ???

OpenStudy (anonymous):

The side lengths are 20 and the width is 8.

OpenStudy (anonymous):

perimeter = l+l+w+w 56 = 2x+2(x+12) 56 = 4x+24 32=4x x=8 Length = 20 Width = 8

OpenStudy (anonymous):

56-12=44 44-12=32 Now you have the extra lengths set apart now you divide. 32/4=8 Your width is 8 8+12=20 your length is 20